# Hierarchical Quantum Control Gates for Functional MRI Understanding

### v2

@@ I Introduction
Brain-inspired representations provide a promising pathway for enhancing the training of deep neural networks (DNN) [1]. To understand these representations, it is crucial to monitor the activities of every single neuron in the brain simultaneously. Among several neurocomputing techniques such as functional Magnetic Resonance Imaging (fMRI), Magnetoencephalography (MEG), and electroencephalogram (EEG), due to the efficiency, it is worth mentioning fMRI signals that have been used widely for that purpose in most recent studies [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. These studies explored patterns of fMRI signals in a trivial approach, i.e., using a fully connected layer to extract the features. Meanwhile, the goal of understanding the pattern of fMRI signals is to discover the relationship between neurons that activate together. For instance, when a human perceives a visual science that contains the face of someone, the voxels within floc-face region inside the brain will be fired or activated. For that reason, the previous approaches fall apart in exploring fMRI patterns.
@@
The limitations when using self-attention to deal with very long sequences are high memory and complexity. Especially, the complexity and memory are quadratic and square of the sequence length. Meanwhile, fMRI signals are extremely high-dimensional signals that contain thousands of voxel activations. Thus, it is still a big obstacle to utilize the self-attention mechanism for understanding fMRI signals.
Apart from classical computers, quantum computers can process information in parallel thanks to qubits’ properties. Specifically, quantum bits, or qubits, can exist in a superposition of states, representing both 0 and 1 simultaneously. It allows a quantum computer to process a vast number of possible outcomes simultaneously, unlike classical bits, which are either 0 or 1. Due to superposition, a quantum computer can evaluate many possible solutions at once. Inspired by this property, we propose a quantum-based solution named a novel approach for fMRI understanding. The Fig. 1 demonstrates the idea of our proposed method. The contributions in this paper can be summarized as follows:
  • We introduce a novel quantum-based solution named Hierarchical Quantum Control Gate (HQCG) for fMRI understanding. This method can learn both local and global information of the signals in parallel.
  • Inspired by controllable gate design, we propose a trainable local circuit to explore the local information of the fMRI signal. Specifically, this module is a self-attention method that automatically learns the voxel-wise correlations.
  • To learn global information, we gather all local information extracted by local circuits and learn pair-wise relationships between this information by using a global circuit.
  • From experimental results, the proposed method performs better than a similar one running on a classical computer. Surprisingly, HQCG also helps to prevent overfitting than classical ones.
@@ II-B Quantum Basics
A quantum bit or qubit is the information carrier in the quantum computing and communication channel. A qubit is a two-dimensional Hilbert space with two orthonormal bases $|0\rangle$ and $|1\rangle$. These computational bases are usually represented as vectors $|0\rangle=[1,0]^{\top}$ and $|1\rangle=[0,1]^{\top}$. Due to the unique qubit characteristic of superposition, the state of a qubit can be represented as the sum of two computational bases weighted by complex amplitudes as $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$, where $\alpha$ and $\beta\in\mathbb{C}$, and $|\alpha|^{2}+|\beta|^{2}=1$. $|\alpha|^{2}$ and $|\beta|^{2}$ are the probability of obtaining states $|0\rangle$ and $|1\rangle$ after multiple measurements, respectively. It gives an advantage in quantum computing over classical computing when the qubits can be entangled. The two qubits $q_{0}$ and $q_{1}$ are entangled when they have a state that cannot be individually represented as a complex scalar times the basis vector.
A quantum state $|\psi\rangle$ can be transformed to another state $|\psi^{\prime}\rangle$ through a quantum circuit represented by a unitary matrix $U$. The quantum state transformation can be mathematically formulated as $|\psi^{\prime}\rangle=U|\psi\rangle$. To get classical information from a quantum state $|\psi^{\prime}\rangle$, quantum measurements are applied by computing the expectation value $H\right>=\langle\psi^{\prime}|H|\psi^{\prime}\rangle$ of a Hermitian matrix $H$.
@@ II-C Parameterized Quantum Circuit
where $L$ is the number of operators. The quantum measurements $H$ are used to retrieve the values of the quantum state for further processing.
PQC uses a hybrid quantum-classical procedure to optimize the trainable parameters iteratively. The popular optimization approaches include gradient descent [29], parameter-shift rule [30, 31], and gradient-free techniques [32, 33]. Despite various PQC training and inference problems [34, 35, 36, 37], the PQC poses a potential approach for deep learning tasks in quantum settings due to its entanglement and superposition properties [38, 39].
@@ III-A Data Encoding
Classical computers operate with bits representing either 0 or 1. In contrast, quantum computers use quantum bits, or qubits, which can exist in superpositions of states, such as $|0\rangle$, $|1\rangle$, or any quantum superposition $\alpha|0\rangle+\beta|1\rangle$. To leverage the computational power of quantum machines, classical data must be transformed into these quantum states.
Currently, there are several approaches for quantum data encoding, such as amplitude, phase, or PQC-based encoding (e.g., $U_{3}$). In this paper, we utilize the amplitude encoding strategy for the following reasons. First, amplitude encoding requires fewer qubits than others. Especially, given an fMRI signal length of $l$ denoted as $v=\left[v_{0},\dots v_{l-1}\right]\in\mathbb{R}^{l}$, the amplitude encoding needs only $\left\lceil\text{log}_{2}(l)\right\rceil$ qubits while phase encoding requires $l$ qubits to encode whole data. Second, the amplitude encoding preserves the natural characteristic of the fMRI signals. In particular, the response of the voxel demonstrates the contribution of this voxel to the processes of the brain. For example, given $i^{th}$ and $j^{th}$ voxel, the expression of $v_{i}>v_{j}$ means the $i^{th}$ voxel is more informative than $j^{th}$ voxel. For that reason, amplitude encoding is the most suitable for encoding the fMRI signal.
@@ V-A Multi-Objects Predictions From fMRI
This task aims to predict which objects a participant perceives based on recorded fMRI signals. The dataset includes fMRI signals from both the left and right hemispheres. Table I illustrates the performance of these individual hemispheres as well as their combination.
@@ V-B Training Stability
In this section, we analyze the training stability of the models on classical and quantum computing for fMRI classification problems. We report the accuracy and AUC for both training and validation during the training process, as shown in Fig 4. Initially, when training with the left or right hemispheres, we observe that the training curves of the classical approach are significantly higher than those of the quantum approach; however, the validation curves for the classical approach are consistently lower than those of the quantum approach. It indicates that the classical approach is more prone to overfitting. Additionally, the classical method’s curves exhibit more fluctuations, while the quantum method’s curves are stable and smooth, demonstrating greater training stability. In conclusion, the quantum approach can prevent overfitting and stabilize the training process, leading to better results than the classical approach. Fig 5 shows the training progress of classical and quantum models on different subjects. Clearly, our approach’s performances and stability consistently perform well on different subjects.

### v3

@@ I Introduction
Brain-inspired representations provide a promising pathway for enhancing the training of deep neural networks (DNN) [1]. To understand these representations, it is crucial to monitor the activities of every single neuron in the brain simultaneously. Among several neurocomputing techniques such as functional Magnetic Resonance Imaging (fMRI), Magnetoencephalography (MEG), and electroencephalogram (EEG), due to the efficiency, it is worth mentioning fMRI signals that have been used widely for that purpose in most recent studies [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. These studies explored patterns of fMRI signals in a trivial approach, i.e., using a fully connected layer to extract the features. Meanwhile, understanding the pattern of fMRI signals aims to discover the relationship between neurons that activate together. For instance, when a human perceives a visual science that contains the face of someone, the voxels within floc-face region inside the brain will be fired or activated. For that reason, the previous approaches fall apart in exploring fMRI patterns.
@@
The limitations when using self-attention to deal with very long sequences are high memory and complexity. Especially, the complexity and memory are quadratic and square of the sequence length. Meanwhile, fMRI signals are extremely high-dimensional signals that contain thousands of voxel activations. Thus, utilizing the self-attention mechanism for understanding fMRI signals is still a big obstacle.
Apart from classical computers, quantum computers can process information in parallel thanks to qubits’ properties. Specifically, quantum bits, or qubits, can exist in a superposition of states, representing both 0 and 1 simultaneously. It allows a quantum computer to process a vast number of possible outcomes simultaneously, unlike classical bits, which are either 0 or 1. Due to superposition, a quantum computer can evaluate many possible solutions at once. Inspired by this property, we propose a quantum-based solution named a novel approach for fMRI understanding. The Fig. 1 demonstrates the idea of our proposed method. The contributions in this paper can be summarized as follows: (1) We introduce a novel quantum-based solution named Hierarchical Quantum Control Gate (HQCG) for fMRI understanding. This method can learn both local and global information of the signals in parallel. (2) Inspired by controllable gate design, we propose a trainable local circuit to explore the local information of the fMRI signal. Specifically, this module is a self-attention method that automatically learns the voxel-wise correlations. (3) To learn global information, we gather all local information extracted by local circuits and learn pair-wise relationships between this information by using a global circuit. (4) From experimental results, the proposed method performs better than a similar one running on a classical computer. Surprisingly, HQCG also helps to prevent overfitting than classical ones.
@@ II-B Quantum Basics
A quantum bit or qubit is the information carrier in the quantum computing and communication channel. A qubit is a two-dimensional Hilbert space with two orthonormal bases $|0\rangle$ and $|1\rangle$. These computational bases are usually represented as vectors $|0\rangle=[1,0]^{\top}$ and $|1\rangle=[0,1]^{\top}$. Due to the unique qubit characteristic of superposition, the state of a qubit can be represented as the sum of two computational bases weighted by complex amplitudes as $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$, where $\alpha$ and $\beta\in\mathbb{C}$, and $|\alpha|^{2}+|\beta|^{2}=1$. $|\alpha|^{2}$ and $|\beta|^{2}$ are the probability of obtaining states $|0\rangle$ and $|1\rangle$ after multiple measurements, respectively. It gives an advantage in quantum computing over classical computing when the qubits can be entangled. The two qubits $q_{0}$ and $q_{1}$ are entangled when they have a state that cannot be individually represented as a complex scalar times the basis vector. A quantum state $|\psi\rangle$ can be transformed to another state $|\psi^{\prime}\rangle$ through a quantum circuit represented by a unitary matrix $U$. The quantum state transformation can be mathematically formulated as $|\psi^{\prime}\rangle=U|\psi\rangle$. To get classical information from a quantum state $|\psi^{\prime}\rangle$, quantum measurements are applied by computing the expectation value $H\right>=\langle\psi^{\prime}|H|\psi^{\prime}\rangle$ of a Hermitian matrix $H$.
@@ II-C Parameterized Quantum Circuit
where $L$ is the number of operators. The quantum measurements $H$ are used to retrieve the values of the quantum state for further processing. PQC uses a hybrid quantum-classical procedure to optimize the trainable parameters iteratively. The popular optimization approaches include gradient descent [29], parameter-shift rule [30, 31], and gradient-free techniques [32, 33]. Despite various PQC training and inference problems [34, 35, 36, 37], the PQC poses a potential approach for deep learning tasks in quantum settings due to its entanglement and superposition properties [38, 39].
@@ III-A Data Encoding
Classical computers operate with bits representing either 0 or 1. In contrast, quantum computers use quantum bits, or qubits, which can exist in superpositions of states, such as $|0\rangle$, $|1\rangle$, or any quantum superposition $\alpha|0\rangle+\beta|1\rangle$. To leverage the computational power of quantum machines, classical data must be transformed into these quantum states. Currently, there are several approaches for quantum data encoding, such as amplitude, phase, or PQC-based encoding (e.g., $U_{3}$). In this paper, we utilize the amplitude encoding strategy for the following reasons. First, amplitude encoding requires fewer qubits than others. Especially, given an fMRI signal length of $l$ denoted as $v=\left[v_{0},\dots v_{l-1}\right]\in\mathbb{R}^{l}$, the amplitude encoding needs only $\left\lceil\text{log}_{2}(l)\right\rceil$ qubits while phase encoding requires $l$ qubits to encode whole data. Second, the amplitude encoding preserves the natural characteristic of the fMRI signals. In particular, the response of the voxel demonstrates the contribution of this voxel to the processes of the brain. For example, given $i^{th}$ and $j^{th}$ voxel, the expression of $v_{i}>v_{j}$ means the $i^{th}$ voxel is more informative than $j^{th}$ voxel. For that reason, amplitude encoding is the most suitable for encoding the fMRI signal.
@@ V-A Multi-Objects Predictions From fMRI
This task aims to predict which objects a participant perceives based on recorded fMRI signals. The dataset includes fMRI signals from both the left and right hemispheres. Table I illustrates the performance of these individual hemispheres and their combination.
@@ V-B Training Stability
In this section, we analyze the training stability of the models on classical and quantum computing for fMRI classification problems. We report the accuracy and AUC for both training and validation during the training process, as shown in Fig 4. Initially, when training with the left or right hemispheres, we observe that the training curves of the classical approach are significantly higher than those of the quantum approach; however, the validation curves for the classical approach are consistently lower than those of the quantum approach. It indicates that the classical approach is more prone to overfitting. Additionally, the classical method’s curves exhibit more fluctuations, while the quantum method’s curves are stable and smooth, demonstrating greater training stability. In conclusion, the quantum approach can prevent overfitting and stabilize the training process, leading to better results than the classical approach. Fig 5 shows the training progress of classical and quantum models on different subjects. Clearly, our approach’s performances and stability consistently perform well on various subjects.
@@ VI Conclusion and Discussion
Acknowledgment. This work is partly supported by MonArk NSF Quantum Foundry, the National Science Foundation Q-AMASE-i program under NSF award No. DMR-1906383, the University of Arkansas Travel Grant, and JBHunt Company.